OpenStudy (anonymous):

substitution method 3x+5y=-3 -7x+y=45

OpenStudy (anonymous):

help with this one pat18 please i'm struggling with these problems substitution method 3x+5y=-3 -7x+y=45

OpenStudy (anonymous):

In these kinds of problems you want to solve for either x or y on ONE of the equations. Let me show you why. The second equation is easier to solve to y so let's do that first. -7x+y=45 Add 7x to both sides. y=45+7x Now you can substitute y into the first equation! So we have 3x+5y=-3, right? Put the value we got for y (45+7x) so: 3x+5(45+x)=-3. Now let's solve it. 3x+225+5x=-3 3x+5x=-228 8x=-228 x=-28.5 Now remember how we solved for the y equation earlier? Now we can substitute the value for x in that equation. y=45+7x. So y=45+7(-28.5). Now let's solve. y=45+7(-28.5) y=45-199.5 y=-154.5 So the solution to this equation is: (-28.5, -154.5)

OpenStudy (anonymous):

It's the solution to the system of equations, rather.*